Consider the following IS-LM model:
C = Co + c1(Y-T)
I = b0 + b1Y -b2i
M/P = d1Y – d2i
a. Solve for equilibrium output. Assume (Assume c1 + b1 < 1).
Now let investment depend on both sales and the interest rate:
b. Solve for equilibrium level of interest rate.
Let’s go behind the scene in the monetary market. Use the equilibrium in the
money market M/P = d1Y – d2i to solve for the equilibrium level of the real money
supply.
How does the real money supply vary with government spending?
1
Expert's answer
2020-04-30T10:10:31-0400

a. Y=C+I

Y= C0 + c1(Y-T) +b0 + b1Y-b2i 

Y=C0+c1Y+c1T+b0 + b1Y-b2i 

Y=C0+c1Y+c1Yd+b0 + b1Y-b2i 

Y-c1Y-b1Y=C0+c1Yd+b0-b2i 

Y(1-c1-b1)=C0+c1Yd+b0-b2i 

b.Y=(1/1-c1-b1+b2d1/d2)(C0+c1T+b0+(b2/d2)M/P)+G)

M/P=Y/V

 d1Y - d2i=Y/V


with increasing government spending, nominal money supply increases, while real money decreases as a result of inflation